NIR OFF-AXIS ICOS SPECTRUM OF THE NITRATE RADICAL : DOES THE VIBRATIONLESS A^{2}E′′ STATE OF NO_{3} UNDERGO STATIC JAHN-TELLER DISTORTION?

\underline{DAVID J. ROBICHAUD}, KANA TAKEMATSU, AND MITCHIO OKUMURA, Arthur Amos Noyes Laboratory of Chemical Physics, California Institute of Technology, Pasadena CA 91125; PIN CHEN AND HERBERT PICKETT, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109; JOHN F. STANTON, Department of Chemistry, University of Texas, Austin, TX, 78712.

The Jahn-Teller effect in the first two excited states of NO_{3} is poorly understood. There is increasing evidence for relatively strong Jahn-Teller and Pseudo-Jahn-Teller couplings in the Ã^{2}E′′ and B̃^{2}E′ states, suggesting that these states cannot be described by conventional Jahn-Teller hamiltonians. Our previous moderate resolution spectra of the forbidden Ã^{2}E′′ ← X̃^{2}A′_{2} transition revealed further evidence of strong Jahn-Teller interactions, but could not establish if the upper state exhibited static Jahn-Teller distortion. The origin band is strictly forbidden by Herzberg-Teller selection rules, and the 0^{0} level can only be accessed via the weak 4_{1}^{0} hot band. Contour analysis of the partially resolved rotational structure could not definitively establish whether the zero-point averaged geometry has D_{3h} or C_{2v} symmetry. We report the first rotationally resolved spectrum of the 4_{1}^{0} vibronic band of NO_{3} recorded by diode laser spectroscopy, using off-axis Integrated Cavity Output Spectroscopy (ICOS).